OpenStudy (anonymous):

what are the steps to solve a quadratic equation? Can you show me with an example?

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\] \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (unklerhaukus):

example \[x^2+2x+1=0\] \[x=\frac{-2\pm\sqrt{2^2-4}}{2}=-1\] \[(x+1)^2=0\]

OpenStudy (anonymous):

Thanks